OpenStudy (anonymous):

The sides and classification of a triangle are given. The length of the longest side is the integer given. What value(s) of x make the triangle?

OpenStudy (anonymous):

1. x, x, 6 acute

Directrix (directrix):

Are you studying the Triangle Inequality Theorem: One side of a triangle is shorter than the sum of the other two?

Directrix (directrix):

1. x, x, 6 --> triangle with 4 sides lol

OpenStudy (anonymous):

Wait how is it 4 sided?

Directrix (directrix):

1, x, x, 6 ?

OpenStudy (anonymous):

1 as in #1 lol

Directrix (directrix):

Look at the attachment. (You did not answer my question about the Triangle Inequality Theorem).

OpenStudy (anonymous):

No. But I'm pretty sure it's on her handout. I just haven't gotten the chance to look through them.

Directrix (directrix):

6 ^2 < x^2 + x^2 if triangle is to be acute. 36 < 2 x^2 18 < x^2

Directrix (directrix):

If we do not know the theorems, we cannot do the problems. The theorems are the map we need.

OpenStudy (anonymous):

Yup, but you get there, so how would you find x?

Directrix (directrix):

Get the theorem list out.

OpenStudy (anonymous):

She numbers her theorems like theorem 7

Directrix (directrix):

We have to solve this before we "get there" without a map or with a map. lol 18 < x^2

OpenStudy (anonymous):

So I don't exactly know if it is the theorem you're talking about

Directrix (directrix):

This is algebra. 18 < x^2 or x^2 >18 x^2 - 18 > 0. Algebra solves that. You could try taking square roots of both sides of x^2 >18.

OpenStudy (kinggeorge):

Since the triangle you're looking for is isosceles and acute, you know that the sum of the squares of the two equal side are greater than the square of the non-equal side. In other words, \[x^2+x^2 > 6^2\]

OpenStudy (kinggeorge):

I thought Directrix summed it up pretty well.

OpenStudy (anonymous):

So couldn't x be 4 or 3?

Directrix (directrix):

x^2 + x^2 < 36, I think.

OpenStudy (kinggeorge):

No, x can't be 4 or 3.

OpenStudy (anonymous):

I get the rest of the steps, just don't understand how to find x.

Directrix (directrix):

Algebra Time: x^2 + x^2 < 36

OpenStudy (kinggeorge):

That should be greater than, not less than.

OpenStudy (kinggeorge):

\(x^2+x^2 > 36\)

OpenStudy (kinggeorge):

So \(2x^2 > 36 \Longrightarrow x^2 > 18\) Since \(36/2 = 18\)

OpenStudy (kinggeorge):

Then, we only have to take the square root of both sides, so \(\sqrt{x^2} > \sqrt{18} \Longrightarrow x > 3\sqrt2\)

OpenStudy (kinggeorge):

Did this make more sense?

OpenStudy (anonymous):

Oh! I see now. I think it was the > sign that threw me off a bit.

OpenStudy (kinggeorge):

It's all good. Do you think you could do a similar problem in the future?

OpenStudy (anonymous):

Yes, in fact, I have one right now. After I finish, will you tell me if I'm correct?

OpenStudy (kinggeorge):

of course